OpenStudy (anonymous):

Solve the following system of equations. x + y – 4z = –7 2x + 3y – 3z = –6 2x + 2y – z = 0

jhonyy9 (jhonyy9):

- subtract 3rd from second 2x +3y -3z = -6 2x +2y -z = 0 --------------- 0 y -2z = -6 - multiply first by -2 and assum by second -2x -2y +8z = 14 2x +3y -3z = -6 ---------------- 0 y +5z = 8 y-2z=-6 y+5z=8 ---------- subtract second from first 0 -7z =-14 z=2 y+5z=8 y+5*2=8 y+10=8 y=-2 z=2 y=-2 x+y -4z=-7 x +(-2) -4(2) = -7 x-2-8=-7 x-10=-7 x=3 x=3 y=-2 z=2

OpenStudy (anonymous):

I hadnt realized you answered thank you so much:D